OpenStudy (anonymous):

Write the equation for a circle centered at the origin with a radius of 14.

OpenStudy (igreen):

Equation of a circle: \(\sf (x - h)^2 + (y - k)^2 = r^2\) Where (h, k) is the center and 'r' is the radius.

OpenStudy (anonymous):

\[(x+1)^{2}+(y-2)^{2}=14^{2}\]

OpenStudy (dmndlife24):

The origin is at (0, 0)

OpenStudy (anonymous):

\[(x-0)^{2}+(y-0)^{2}=14^{2} \]

OpenStudy (dmndlife24):

Correct